A sample is selected from a population with M=80.
If the sample has a mean of 85,which of the following will produce the largest value for t-obtained ?

A)n=16 and s^2=16
B)n=36 and s^2=36
C)n=16 and s^2=36
D)n=36 and s^2 =16
Answer #1

A)

The test statistics:

M_μ 85-80 M-11

B)

The test statistics:

t=\frac{M-\mu}{s/\sqrt{n}}=\frac{85-80}{\sqrt{36}/\sqrt{36}}=5

C)

The test statistics:

t=\frac{M-\mu}{s/\sqrt{n}}=\frac{85-80}{\sqrt{36}/\sqrt{16}}=3.333

D)

The test statistics:

t=\frac{M-\mu}{s/\sqrt{n}}=\frac{85-80}{\sqrt{16}/\sqrt{36}}=7.500

Correct option is D.